OpenStudy (anonymous):

The formula for the volume of a sphere is y = 4/3 pi r^3 . a. Solve the formula for r and write using positive rational exponents. b. Find the radius, to the nearest hundredth, of a sphere with a volume of 15 in.3

OpenStudy (phi):

You posted this question yesterday. Do you need more help? Solve the formula for r and write using positive rational exponents. can you solve for r^3 (get r^3 by itself on side of the = sign) ?

OpenStudy (phi):

for info on using "rational exponents" (fancy word for a fraction used as an exponent) see http://www.khanacademy.org/math/arithmetic/exponents-radicals/world-of-exponents/v/zero--negative--and-fractional-exponents

OpenStudy (anonymous):

so is this right for the first part of the problem? \[r = \sqrt[3]{\frac{ 3y }{ pi }}\]

OpenStudy (phi):

almost, you left out the 4 \[r=\sqrt[3]{\frac{3y}{4 \pi} }\]

OpenStudy (phi):

now use this idea \[ \sqrt[n]{x} = x^{\frac{1}{n}}\]

OpenStudy (anonymous):

im confused now

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Did you look at the video?

OpenStudy (phi):

y = 4/3 pi r^3 \[ y = \frac{4}{3} \pi r^3\] In little steps: multiply both sides by 3 \[ 3y = \cancel{3}\cdot \frac{4}{\cancel{3}} \pi r^3 \] divide both sides by 4π \[ \frac{3y}{4\pi} = \frac{\cancel{4\pi}}{\cancel{4\pi}} r^3\] \[ r^3 = \frac{3y}{4\pi} \] now raise both sides to the 1/3 power \[ \left(r^3\right)^{\frac{1}{3}} = \left(\frac{3y}{4\pi}\right)^{\frac{1}{3}} \] you should know (or watch more videos!) that \[ (x^a)^b = x^{ab} \] for this problem x is r, a is 3 and b is 1/3, and we can re-write the left side as \[ r^{3\cdot \frac{1}{3}} = r^1= r\] your answer is \[ r= \left(\frac{3y}{4\pi}\right)^{\frac{1}{3}} \]

OpenStudy (anonymous):

and then you plug in 15 for y right

OpenStudy (phi):

yes, you can type into google (45/(4*pi))^(1/3) = or cube root (45/(4*pi)) =

OpenStudy (anonymous):

thank you so much